### v1

# Small Open Models Achieve Near-Parity with Large Models in Low-Resource Literary Translation at a Fraction of the Cost

Abstract
[…]
1 Introduction
Translating literature into low-resource languages poses unique challenges, given both the scarcity of parallel data and the need to faithfully preserve stylistic nuance and cultural context. Romanian—a Latin language spoken by over 24 million people—remains notably underserved in the realm of high-quality literary translation resources. Existing machine translation (MT) benchmarks, such as WMT [^1] WMT stands for the Conference on Machine Translation (formerly “Workshop on Machine Translation”). The WMT organizes annual shared tasks for benchmarking machine translation systems, with its “news translation” task being the most widely used standard evaluation for translation of contemporary news articles. See: https://www.statmt.org/wmt23/., predominantly focus on informational or news text, thus providing little guidance for creative, narrative-driven translation tasks. The rapid advancement of large language models (LLMs) has, however, enabled synthetic text generation at unprecedented scale, opening up new opportunities for augmenting MT datasets [14, 19]. Notably, projects like TinyStories [4] have shown that even compact models can produce coherent narratives when trained on thoughtfully curated synthetic corpora. While these advances have transformed data generation and model pretraining, much of the prior work has centered on monolingual story generation in English [6, 4] or improvements to translation via back-translation [22], rather than the systematic creation of large-scale, bilingual literary datasets. As a result, there remain open questions about how both proprietary and open-source LLMs can be harnessed to create and evaluate high-quality literary translations in genuinely low-resource settings.
Significant progress has recently been made toward building massive bilingual and multilingual resources—for example, CLIRMatrix [26] compiles parallel data across 139 languages for cross-lingual retrieval, and mT5 [33] pretrains a text-to-text transformer on a vast array of language pairs and domains. However, such resources are primarily oriented toward general-domain or retrieval-centric MT, with limited representation of literary content. Even in language pairs like English–Romanian, coverage is often restricted to news, web, or miscellaneous domains. To date, there is no open, large-scale, high-quality English–Romanian parallel corpus specifically curated for creative narratives or fables.
TinyFabulist Translation Framework (TF2) addresses this gap. DS-TF2-EN-RO-3M is a 3-million-example synthetic dataset of English–Romanian parallel fables, accompanied by fine-tuned open-source translation models and a literary translation benchmark. It extends our earlier TinyFabulist-TF1 corpus of English-only fables [18] into a bilingual setting, explicitly targeting creative narrative translation and cultural adaptation. By combining parameter-efficient adaptation with narrative-aware evaluation, TF2 provides both a large-scale resource and a practical benchmark for studying literary MT in a genuinely low-resource language pair.
While recent multilingual models such as NLLB-200 [27] and EuroLLM [16] have demonstrated impressive performance, these systems are generally optimized for breadth and general-domain translation. Their large parameter counts also limit practical deployment in cost-constrained or on-device scenarios. Moreover, our experiments show that such models, despite their coverage, often fall short in producing the nuanced style, coherence, and cultural adaptation required for professional literary translation (see Section 4).
Despite recent advances in both synthetic data creation and multilingual modeling, there remains a critical lack of resources and benchmarks for literary translation in low-resource settings—especially for genres demanding nuanced handling of narrative style and cultural meaning. Existing open and proprietary translation models are typically not tuned for creative content, and English–Romanian parallel datasets remain scarce, limited in scope, or non-public. Prior benchmarks overwhelmingly favor news or technical content, leaving literary and didactic narratives comparatively neglected.
Building on the English-only fable generation efforts of DS-TF1-EN-3M [18], we systematically extend to cross-lingual translation, focusing on moral fables—short stories that require not only semantic accuracy but also fluency, narrative coherence, stylistic faithfulness, and nuanced cultural adaptation. Generating high-quality parallel data under resource constraints is especially challenging for such genres, motivating our use of LLMs for both dataset creation and the development of an automated, multi-dimensional evaluation framework.
To address the challenges outlined above, our work makes the following key contributions:
  • Narrative-Aware Evaluation Protocol. We show how to bootstrap high-quality literary data with instruction-tuned LLMs, combining corpus-level BLEU [25] with a five-dimension LLM rubric for reproducible, fine-grained assessment.
[…]
  • Three open, fine-tuned translation models. We release three open, fine-tuned translation models—tf2-1b, tf2-4b, and tf2-12b—LoRA-adapted checkpoints that substantially close the quality gap to proprietary systems while remaining efficient enough to run locally on commodity GPUs. [^3] https://huggingface.co/klusai/tf2-1b,  https://huggingface.co/klusai/tf2-4b,  https://huggingface.co/klusai/tf2-12b
  • Fully transparent artifacts. All data, code, prompts, and evaluation scripts are published under permissive licenses to foster reproducibility and further research in low-resource literary NLP.
By focusing on literary translation in a genuinely low-resource language, TF2 fills a gap left by existing datasets and MT models, demonstrating that high-quality, culturally adapted translation is feasible within tight cost and resource constraints.
Building on this foundation, the remainder of the paper is organized to systematically explore the practical and methodological questions that arise in the creation and evaluation of large-scale literary translation resources for low-resource languages. We motivate our experimental design, comparative analyses, and evaluation framework in the context of both real-world constraints and the broader landscape of open and proprietary language technologies.
1.1 Research Questions
In formulating our research questions, we recognize that the development of large-scale translation datasets and models is subject to a variety of practical constraints. Among these, financial cost remains the most immediate and stringent limitation, especially for low-resource language communities or research teams with restricted budgets. However, other important factors must also be considered, including the environmental impact associated with large-scale computation—such as CO2 emissions and water use for hardware cooling—as well as the ethical provenance of training data, particularly with respect to data privacy and fair representation. While our primary focus in this work is on optimizing inference for a domain-specific NLP task, we acknowledge that by doing so we also address sustainable and responsible machine translation research, hence helping to reduce ecological footprint and adhere to ethical data standards. These broader considerations motivate the structure of our experimental pipeline and inform our discussion of the trade-offs inherent in large-scale synthetic data generation.
Accordingly, we structure our study around the following central research questions (RQs):
  1. RQ1
    Cost-Constrained Data Generation:
  2. RQ1
    Cost-Constrained Data Generation: Can a large-scale, high-quality English–Romanian literary translation dataset (moral fables) be built under strict budget constraints—orders of magnitude below state-of-the-art (SOTA) proprietary LLMs—using open models of 12B parameters or fewer, and what trade-offs are necessary to achieve this?
  3. RQ2
    Open vs. Proprietary Translation Quality: How do SOTA proprietary models (such as GPT-o3 and Google’s Gemini) compare to open-source models (before and after fine-tuning) in terms of Romanian literary translation quality? Can a compact open model, fine-tuned on the generated fable corpus, approach the translation fidelity and fluency of much larger proprietary systems?
  4. RQ3
    Evaluation Rigor and Automation: Can an automated evaluation framework using an LLM-as-a-judge reliably assess literary translation quality across multiple dimensions (accuracy, fluency, coherence, style, cultural adaptation)? How well do these LLM-based evaluations align with expectations of professional literary translation, and what are the limitations of this approach?
By addressing RQ1, we aim to establish a practical pipeline for low-cost dataset creation. RQ2 examines the effectiveness of budget-constrained open models relative to the best available closed models, shedding light on the trade-offs between accessibility and performance. RQ3 focuses on ensuring that our evaluation methodology is robust and can capture the nuances of literary translation better than crude metrics like BLEU alone.
2 Related Work
Synthetic data and low-resource Machine Translation (MT).
Large language models have been widely used to fabricate training corpora for tasks where human data are scarce or costly. Early work on back-translation [22] showed that automatically generated target–source pairs can boost neural MT in low-resource regimes. More recently, LLM-driven pipelines have produced massive monolingual or parallel datasets at a fraction of traditional annotation cost [14, 31]. In narrative domains, Eldan and Li [4] created TinyStories, demonstrating that millions of synthetic tales enable coherent story generation with sub-10M-parameter models. Our TF2 corpus follows this line of research but focuses on bilingual literary content—an area where synthetic data remain under-explored.
Fable generation and moral reasoning.
Moral or didactic stories have attracted interest as test-beds for value-aligned generation and commonsense reasoning. Guan et al. [6] introduced STORAL, a human-authored dataset coupling narratives with morals; subsequent work explored moral inference and story completion. TinyFabulist DS-TF1-EN-3M [18] scaled synthetic moral fables to three million English examples. TF2 extends this effort cross-lingually, supplying the first large-scale English–Romanian parallel fable corpus and highlighting translation-specific challenges such as cultural adaptation of morals.
Open vs. proprietary LLMs for translation.
Comparative studies of closed APIs (e.g., GPT-4, Gemini) and open-weight models (Llama, Mistral, Qwen) report a measurable but narrowing quality gap, especially for less-represented languages [36]. Parameter-efficient fine-tuning methods—most notably LoRA adapters [7] and subsequent PEFT toolkits [32]—have demonstrated substantial gains for open models, closing much of the gap on task-specific benchmarks while maintaining low deployment costs. Recent open-weight multilingual giants such as DeepSeek-LLM [1], EuroLLM-9B [16], and Meta’s NLLB-200 [27] achieve impressive translation quality across dozens to hundreds of languages, including Romanian. However, their large parameter counts and memory requirements place them out of reach for many budget- or hardware-constrained users. By contrast, TF2 focuses on parameter-efficient adaptation of compact and mid-sized open models (1B, 4B, and 12B parameters), demonstrating that with careful data curation, instruction tuning, and adapter compression, even relatively lightweight models can approach the translation quality of much larger proprietary systems for creative literary tasks. This three-pronged strategy supports scalable, cost-effective deployment, while still yielding strong empirical results in both automated and rubric-based evaluation.
Evaluation beyond BLEU.
While BLEU [20] remains the de-facto MT benchmark, its correlation with human judgment drops for creative text. Frameworks such as MQM [13] and COMET [21] address adequacy and fluency more directly, but require either expert annotators or supervised reference models. Recent work leverages LLMs as automatic judges, achieving human-level agreement for summarisation and dialogue evaluation [12]. We adopt this LLM-as-judge paradigm, extending it with a five-dimension rubric tailored to literary translation—accuracy, fluency, coherence, style, and cultural adaptation.
Several recent approaches push evaluation further in ways highly relevant to narrative translation. Hu et al. [8] propose Cont-COMET, a context-aware extension of COMET that incorporates preceding and following sentences to better align with human annotations, improving both system-level and segment-level assessments. This context-sensitive design highlights how broader narrative coherence could also be integrated into evaluation—a direction we aim to extend in future work. Wang et al. [29] introduce a multi-agent system where different LLMs assess terminology consistency, narrative perspective, and stylistic fidelity, combining their judgments into an overall translation quality score. Zhang et al. [35] present LiTransProQA, a QA-based evaluation framework where LLMs answer targeted questions about style, tone, and cultural references, mapping responses into quantitative scores. Finally, Yao et al. [34] release the CAMT corpus alongside metrics focused on culture-specific items, showing that LLM-based evaluation outperforms neural MTs in cultural adaptation.
Cost-aware NLP pipelines.
Finally, a growing body of work emphasises budget-constrained model development, from efficient inference engines [5] to dataset generation strategies that trade minimal API spend for maximal downstream benefit [30]. TF2 contributes an end-to-end case study: dataset synthesis, fine-tuning, and evaluation—all within a moderate budget by current standards—thus offering a practical blueprint for researchers with limited compute or funding.
Open-source Romanian language resources.
Recent efforts have started to address the scarcity of open benchmarks and models for Romanian natural language processing. Masala et al. [17] present OpenLLM-Ro, a comprehensive suite of open-source Romanian language models, evaluation benchmarks, and datasets, released through both GitHub [^4] https://github.com/openllm-ro and Hugging Face [^5] https://huggingface.co/OpenLLM-Ro. These resources provide a foundation for research in Romanian LLMs, but, to our knowledge, do not specifically target literary translation or creative narrative domains. Our TF2 benchmark is complementary, focusing on literary translation tasks and on the cross-lingual transfer of narrative content, thereby addressing a different—yet crucial—aspect of Romanian language technology.
Taken together, these strands of research underscore the feasibility and importance of cost-efficient, open-model pipelines for literary translation in low-resource languages. Our work contributes to this growing field by introducing TF2 based on an approach that combines synthetic data generation, parameter-efficient fine-tuning, and narrative-aware evaluation to produce high-quality English–Romanian literary translations under strict budget constraints.
3 TinyFabulist Translation Framework (TF2)
TF2 generalizes the original TinyFabulist Framework (TF) project to a bilingual and scalable setting, introducing an end-to-end methodology for constructing a large-scale, high-quality English–Romanian literary translation resource. The framework is architected for transparency, replicability, and practical cost-awareness at every step, emphasizing open data, parameter-efficient adaptation, and rigorous evaluation.
The workflow (see Figure 1) consists of the following four main stages, each described and justified in detail below:
  1. S1
    Evaluating Candidate Translators: We benchmark 13 diverse LLMs and commercial translation APIs on a held-out fable set (e.g. in our case the DS-TF1-EN-3M dataset [18]), with evaluation targeting not only accuracy but also literary style and coherence. Metrics include a five-dimension LLM-based rubric, capturing the specific requirements of creative translation.
  2. S2
    Parallel Dataset Creation: The best-performing Stage 1 system is used to automatically translate 15.000 English fables, producing the TinyFabulist DS-TF2-EN-RO-15K corpus. This “silver-standard” parallel set supplies essential supervision for downstream model adaptation in a domain lacking human references, and also introduces a ground-truth reference enabling BLEU-based benchmarking.
  3. S3
    Parameter-Efficient Fine-Tuning: A suite of open LLMs—ranging from 1B to 12B parameters—are fine-tuned on the 15k parallel set using Low-Rank Adaptation (LoRA [7]), yielding domain-specialized English–Romanian translation models. Multiple quantized variants are produced for efficient local inference. The fine-tuning process is designed to be transparent and reproducible.
  4. S4
    Large-Scale Corpus Generation: The best fine-tuned model(s) are used to translate the remaining $\sim$3M English fables from DS-TF1-EN-3M dataset, resulting in the largest openly available bilingual literary dataset for Romanian. The pipeline is modular, enabling ongoing evaluation and future expansion as models or translation strategies improve.
Each stage is tightly integrated with rigorous quality control and open resource release, providing a replicable blueprint for low-resource literary MT beyond English–Romanian.
[…]
Figure 1: TinyFabulist Translation Framework (TF2) pipeline: Evaluation of translation models on literary benchmarks (S1); creation of a 15k English–Romanian parallel corpus via best system (S2); parameter-efficient fine-tuning of open LLMs and quantized variants (S3); and large-scale Romanian translation of the full English corpus (S4).
3.1 Stage 1: Evaluating Models & Selecting Translators
The first stage identifies (i) a high-quality translation system to serve as a reference translator and (ii) a competitive open-source backbone suitable for fine-tuning. To this end, we evaluated a diverse pool of candidate systems, including commercial MT APIs, proprietary LLMs, and open-source instruction-tuned models, to gauge their suitability for the literary translation domain.
Evaluation Methodology. Because standard automatic metrics such as BLEU are poorly correlated with literary adequacy and stylistic fidelity, we relied exclusively on LLM-based human-like evaluation. Each candidate was scored on five dimensions derived from professional translation guidelines:
  1. 1.
    Accuracy: Fidelity to the full semantic content of the source.
  2. 2.
    Fluency: Grammaticality, naturalness, and readability in the target language.
  3. 3.
    Coherence: Logical flow and clarity at the story level.
  4. 4.
    Style: Appropriateness of tone and consistency with the narrative voice.
  5. 5.
    Cultural/Pragmatic Adaptation: Adequacy of cultural references and moral framing for the target audience.
Each dimension is scored on a 1 to 5 scale (1 = poor, 5 = excellent). To conduct this evaluation at scale, we utilize an LLM as an automated judge. This model was prompted with detailed instructions to emulate a professional translation reviewer. The prompt (in natural language) set the context: “You are a professional translation evaluator. You will be given an English fable and a Romanian translation. Evaluate the translation for accuracy, fluency, coherence, style, and cultural/pragmatic fidelity. Provide a score from 1 to 5 for each category along with a brief justification. Output your evaluation in valid JSON format with fields for each score and justification.”
By constraining the output to JSON, we ensured the evaluator’s responses could be easily parsed and aggregated. We evaluated a randomly selected subset of 100 test instances with this LLM-based evaluator for each model’s translations (100 unique fables per model). The result is a set of scores across the five dimensions, as well as qualitative explanations (which we used to sanity-check the scores). Using an LLM as a judge in this manner draws on recent work such as G-Eval and GPTScore, which have shown that LLMs can approximate human evaluation of text generation to a useful degree [12]. The average scores from these evaluations are reported in Table 1, providing a multi-dimensional comparison of translation quality.
All evaluation prompts, along with example JSON outputs, are included in our repository for transparency. The JSON output for each evaluated sample contains keys like "accuracy": 5, "fluency": 4, and so forth. We emphasize that these LLM-generated scores are used as an auxiliary signal, not as an absolute ground truth; nonetheless, they offer valuable insight into where certain models might be failing (e.g., lower style scores despite strong accuracy).
Selecting Translators and Backbones. From these evaluations, the system with the highest rubric scores was designated as the reference translator for corpus creation and benchmarking. Separately, the strongest-performing open-source model was selected as the backbone for fine-tuning in Stage 3. This dual selection ensures both high-quality references for evaluation and open, reproducible foundations for model adaptation.
3.2 Stage 2: Parallel Corpus Generation (15k Fables)
A central step in constructing a robust benchmark for low-resource literary MT is the creation of a high-quality parallel dataset tailored to the narrative and stylistic demands of the target genre. In TF2, we address the acute shortage of English–Romanian literary data by leveraging state-of-the-art LLMs to synthesize a sizable and diverse set of story translations, adhering to the best practices in synthetic corpus construction [14, 4].
The pipeline begins by sampling 15,000 short fables from the DS-TF1-EN-3M dataset [^6] https://huggingface.co/datasets/klusai/ds-tf1-en-3m, each consisting of a concise narrative culminating in an explicit moral. Rather than relying on scarce human-translated sources, we use the top-performing LLM-based translator, selected via rigorous benchmarking, to generate Romanian counterparts for each English story. This process yields a high-coverage, machine-generated parallel corpus, designed to serve as both a training resource and a reference benchmark for literary MT.
The resulting TinyFabulist DS-TF2-EN-RO-15K dataset is released on Hugging Face [^7] https://huggingface.co/datasets/klusai/tf2-en-ro-15k under the permissive MIT License, ensuring full transparency and broad community reuse. Each record is stored in JSONL (one JSON object per line) and includes the English source, the Romanian target, and comprehensive metadata to facilitate downstream analysis and reproducibility. Key fields are:
  • fable: The original English fable text, typically 1–3 paragraphs, always ending with an explicit moral.
  • translated_fable: The Romanian translation generated by the selected model.
  • pipeline_stage: Stage of the data-generation pipeline that produced the record (e.g., translation).
  • source_lang and target_lang: Language codes indicating source and target (here always English $\rightarrow$ Romanian).
  • prompt_hash: SHA-256 hash of the generation prompt, used for deduplication and reproducibility.
  • llm_name: Identifier of the model used for translation (path or Hugging Face snapshot).
  • translation_model: Duplicate field pointing to the model checkpoint that produced the translation, for explicit traceability.
  • generation_timestamp: Unix timestamp of when the translation was generated, enabling chronological auditing.
The dataset is split into 12,000 training pairs, 1,500 validation pairs, and 1,500 test pairs, supporting robust supervised learning and benchmarking. Although the Romanian references are model-generated, their quality is ensured by selecting the strongest available translator through the quality-aware benchmarking pipeline outlined in Section 3.1. This pragmatic approach is aligned with recent best practices for corpus creation in low-resource settings, and each entry’s rich metadata supports reproducibility, cross-evaluation, and longitudinal studies.
Overall, the TinyFabulist DS-TF2-EN-RO-15K corpus provides a transparent, extensible foundation for training, evaluation, and methodological innovation in literary machine translation, directly addressing the lack of large-scale, genre-specific resources for English–Romanian and similar language pairs.
3.3 Stage 3: Fine-Tuning—Parameter-Efficient Domain Adaptation
The third stage adapts open instruction-tuned LLMs to the literary translation domain using parameter-efficient fine-tuning. Instead of updating all model parameters, which is costly and less reproducible, we employ Low-Rank Adaptation (LoRA; 7), a technique that injects lightweight, trainable adapters into attention and feed-forward layers while keeping most parameters frozen. This approach yields strong domain adaptation with minimal computational overhead.
Training Methodology. The fine-tuning process follows established instruction-tuning practices:
  1. (a)
    Data preparation: Align and preprocess English–Romanian text pairs into instruction–response format, filtering for maximum length and domain relevance. Each input is prefixed with a standardized prompt (e.g., “Translate the following fable from English to Romanian:”).
  2. (b)
    Tokenization and masking: Employ the native SentencePiece or BPE vocabulary associated with each model family. In the labels tensor, mask all tokens except the Romanian output, ensuring that only target-side predictions contribute to the training loss. English inputs are tokenized but ignored in the loss function; only target tokens contribute to gradients.
  3. (c)
    Adapter injection: We attach LoRA adapters to all major projection matrices in both self-attention and feed-forward blocks—q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. For each weight $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ we learn a low-rank update $\Delta W=\tfrac{\alpha}{r}BA$ with r=32r{=}32, $\alpha{=}32$ (effective scaling =1.0=1.0), and apply dropout p=0.05p{=}0.05 on the adapter path. We train only $A,B$ while freezing the backbone (no bias adaptation), and merge $\Delta W$ into $W$ at inference. This follows best practices in recent literature [15, 3].
  4. (d)
    Training regime: Fine-tune models using AdamW optimization, gradient accumulation, cosine learning rate scheduling, and mixed-precision (FP16 or bfloat16) arithmetic. To maximize sample efficiency and prevent overfitting, we combine LoRA dropout with early stopping based on held-out validation loss, following Liu et al. [12].
  5. (e)
    Deployment: After training convergence, we merge LoRA adapters into the base weights for standalone inference and produce 8-bit quantized variants [2, 5] to further reduce memory and latency overhead. We additionally leverage llmcompressor to generate W8A8 quantized models, which integrate seamlessly with vLLM for efficient inference [10, 23]. For offline distribution we also export GGUF artifacts compatible with llama.cpp and Hugging Face GGUF endpoints888 https://github.com/ggml-org/llama.cpp.
Evaluation. Fine-tuned models are assessed using both rubric-based evaluation (as in Stage 1) and BLEU. Here, BLEU is used as a complementary metric by comparing translations against the reference corpus created in Stage 2. While rubric scores remain the primary measure of literary adequacy, BLEU provides a lightweight consistency check on lexical overlap, complementing rubric scores without serving as a stand-alone measure.
Backbone Coverage. This fine-tuning procedure is applied across multiple model scales, from compact to large, to demonstrate robustness of domain adaptation under varying computational constraints. Results are reported in Section 4, highlighting improvements in both rubric scores and BLEU relative to untuned backbones.
3.4 Stage 4: Large-Scale Fable Translation
The final stage of the framework is the large-scale translation of the entire TinyFabulist English corpus into Romanian, leveraging our fine-tuned models for efficient and high-quality generation. DS-TF1-EN-3M comprises approximately 3 million AI-generated English fables, all of which were systematically translated into Romanian using our best-performing system, as well as other fine-tuned checkpoints where appropriate. This process produced the DS-TF2-EN-RO-3M dataset: a massive, openly available English–Romanian parallel corpus specifically tailored for literary translation research.
Thanks to the efficiency of the TF2 models, the translation of millions of fables was completed on modest hardware (e.g., small GPU clusters or standard cloud infrastructure) without prohibitive computational cost. The resulting dataset, which vastly surpasses the initial 15k seed corpus in both size and diversity, is now released for the research community and accessible on the Hugging Face Hub. [^9] https://huggingface.co/datasets/klusai/tf2-en-ro-3m
A full statistical and structural overview of the DS-TF2-EN-RO-3M dataset—including metadata schema, field definitions, and example use cases—is provided in Section 5. This large-scale resource supports new directions in literary MT, cross-lingual narrative analysis, and low-resource NLP, furthering the “living corpus” philosophy of TinyFabulist.
Complete hardware and software specifications are listed in Appendix Appendix A. Cost Calculation and Hardware Configurations.
4 Experiments and Results
This section evaluates TF2 ’s translation models against a wide array of strong baselines—both proprietary and open—using both automated and LLM-based metrics. We present a comprehensive comparison of all fine-tuned KlusAI models, their quantized variants, and the influence of decoding temperature on translation quality.
4.1 Benchmarked Systems and Baselines
We first benchmark only external systems as obtained “out of the box” from their public endpoints or hubs, with no additional fine-tuning. The pool spans:
  • Proprietary LLMs and commercial MT: GPT-4.1, GPT-4.1-mini, GPT-o3, GPT-o3-mini, Gemini-2.5, Gemini-2.0, Grok-3, DeepL.
  • Open instruction-tuned baselines: EuroLLM-9B, Qwen3-14B, and untuned Gemma-3 models (1B/4B/12B).
For each model we compute the per-dimension rubric and the overall average:
$\text{Avg.\ Score}=\frac{\text{Accuracy}+\text{Fluency}+\text{Coherence}+\text{Style}+\text{Cultural/Pragmatic}}{5}.$
Model Accuracy Fluency Coherence Style Cultural Avg. Score Count
GPT-o3 (2025-04-16) 4.86 4.92 4.89 4.96 4.97 4.92 100
GPT-4.1-mini (2025-04-14) 4.54 4.71 4.72 4.84 4.83 4.73 98
GPT-4.1 (2025-04-14) 4.86 4.89 4.85 4.92 4.94 4.89 100
GPT-o3-mini (2025-01-31) 4.71 4.78 4.87 4.85 4.92 4.83 100
Gemini-2.5-Flash 4.75 4.86 4.82 4.87 4.89 4.84 100
Gemini-2.0-Flash-001 4.66 4.82 4.78 4.89 4.93 4.82 100
Gemini-Flash-1.5-8b 4.14 4.45 4.67 4.52 4.46 4.45 99
DeepL 4.42 4.73 4.38 4.69 4.74 4.59 100
Grok-3-mini-beta 4.73 4.74 4.77 4.82 4.88 4.79 100
EuroLLM-9B-Instruct 3.84 4.27 4.36 4.27 4.22 4.19 98
Gemma-3-12B-it 3.98 4.56 4.65 4.52 4.43 4.43 100
Gemma-3-4B-it 3.27 3.94 4.17 3.91 3.78 3.81 100
Gemma-3-1B-it 1.79 2.13 2.23 2.07 1.86 2.02 100
Table 1: Non-TF2 baselines on the TF2 literary translation benchmark. Columns are rubric scores (max 5), their mean (Avg. Score), and sample count. Bold indicates the best value per rubric column. Each input sequence contains on average 300–400 tokens, with outputs averaging 350–450 tokens.
Why GPT-o3 as the reference translator?
Table 1 shows that GPT-o3 attains near-ceiling values across all rubric dimensions (4.86–4.97), surpassing the next-best systems (e.g., GPT-4.1, Gemini-2.5) by a consistent margin. We therefore adopt GPT-o3 outputs as our silver-standard references for automatic comparisons elsewhere in the paper. Importantly, subsequent BLEU numbers reported for other systems use these GPT-o3 translations as references, which explains why GPT-o3 itself does not appear with a comparable BLEU entry in those tables.
Why Gemma-3-12B-it as the open backbone to fine-tune?
Among open baselines, Gemma-3-12B-it is the strongest model on our rubric (Avg. 4.43), clearly outperforming the 4B (3.81) and 1B (2.02) variants. It also provides a permissive license and robust tooling, making it a practical and competitive foundation for adaptation. Consequently, we select Gemma-3-12B-it (and, for scaling experiments, its 4B and 1B counterparts) as the backbones for the TF2 fine-tuning in the next subsection.
4.2 Overall Translation Quality with TF2 Models
TF2 fine-tuned models.
The TF2 series comprises three parameter-efficiently fine-tuned variants of the Gemma-3 instruction-tuned backbones: Gemma-3-1B-it, Gemma-3-4B-it, and Gemma-3-12B-it. Each model was adapted on the DS-TF2-EN-RO-15K corpus using LoRA adapters injected into all major projection matrices (q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj) with configuration r=32r{=}32, $\alpha{=}32$, and dropout p=0.05p{=}0.05. Training followed standard low-rank adaptation practice [15], with adapters merged into the base weights after convergence. For deployment, we additionally provide 8-bit quantized variants, as well as a distilled TF2-1B checkpoint derived from teacher–student compression. This setup yields three scalable, self-hostable translation models—TF2-1B, TF2-4B, and TF2-12B—covering the small, medium, and large open-model parameter ranges. Building upon these strong, diverse baselines, we next evaluate all three fine-tuned TF2 models—TF2-1B, TF2-4B, TF2-12B—each directly compared to its original backbone: Gemma-3-1B-it, Gemma-3-4B-it, and Gemma-3-12B-it, respectively. This strategy enables a direct and transparent assessment of the gains achieved through parameter-efficient adaptation on domain-specific, synthetic literary data.
All three TF2 models show dramatic improvements over their untuned checkpoints, closing much of the gap to the top-tier proprietary LLMs. To probe decoding robustness, we evaluate each model under different temperature ($T$) settings: $T=0.0$ (deterministic greedy decoding), $T=0.2$ (low-variance sampling), and $T=1.0$ (higher-variance sampling). Below, we summarize the key results for each parameter scale.
  • TF2-1B (Gemma-3-1B-it backbone): The smallest model, after fine-tuning, attains an average rubric score of 3.75, representing a significant jump from its untuned base (2.02). BLEU, however, remains modest (0.2180 for the TF2-1B-distilled variant; 0.0543 for the standard TF2-1B checkpoint), reflecting the inherent limitations of compact architectures for long-form narrative translation.
  • TF2-4B (Gemma-3-4B-it backbone): The 4B parameter model (both FP16 and 8-bit quantized) consistently achieves averages above 4.65 across all five evaluation dimensions, peaking at 4.74 (T=0.0). BLEU ranges from 0.0496 (quantized, T=0.2) to 0.1290 (quantized, T=0.0) and 0.1278 (FP16, T=1.0), highlighting strong quantization and decoding robustness. This represents a substantial qualitative improvement over the untuned Gemma-3-4B-it base (avg. rubric 3.81, BLEU 0.1005).
  • TF2-12B (Gemma-3-12B-it backbone): The largest model in our suite delivers the strongest results among all open, self-hostable systems. At optimal decoding ($T=0.0$), TF2-12B attains an average rubric score of 4.83 and BLEU 0.0926; at $T=0.2$, 4.82 and BLEU 0.0647; even at $T=1.0$, it maintains a robust 4.66 (BLEU 0.0784). The 8-bit quantized version is nearly indistinguishable, , in line with prior findings on activation-aware quantization [11], showing that aggressive quantization incurs negligible quality loss. All these numbers represent clear, multi-dimensional gains over the untuned Gemma-3-12B-it base (avg. rubric 4.43, BLEU 0.0214).
A summary of main results—with both BLEU and LLM-based rubric scores—is provided in Tables 2 and 3. For each model scale, we report the fine-tuned TF2 model, its quantized variant (where applicable), and the corresponding Gemma backbone for direct comparison.
Model BLEU Score Notes
TF2-12B (T=0.0) 0.0926 fine-tuned (FP16)
TF2-12B (T=0.2) 0.0647 fine-tuned (FP16)
TF2-12B (T=1.0) 0.0784 fine-tuned (FP16)
TF2-12B-quant (T=0.0) 0.0746 quantized (8-bit)
TF2-12B-quant (T=0.2) 0.0644 quantized (8-bit)
TF2-12B-quant (T=1.0) 0.0548 quantized (8-bit)
Gemma-3-12B-it 0.0214 untuned base
TF2-4B (T=0.0) 0.1153 fine-tuned (FP16)
TF2-4B (T=0.2) 0.1094 fine-tuned (FP16)
TF2-4B (T=1.0) 0.1278 fine-tuned (FP16)
TF2-4B-quant (T=0.0) 0.1290 quantized (8-bit)
TF2-4B-quant (T=0.2) 0.0496 quantized (8-bit)
TF2-4B-quant (T=1.0) 0.0857 quantized (8-bit)
Gemma-3-4B-it 0.1005 untuned base
TF2-1B 0.0543 fine-tuned (T=0.5)
TF2-1B-distilled 0.2180 distilled
Gemma-3-1B-it 0.0790 untuned base
Table 2: BLEU scores for each fine-tuned TF2 model and its corresponding Gemma backbone. Quantized and distilled variants are included. BLEU reflects test set n-gram overlap with the GPT-o3 reference and is reported on a normalized 011 scale ($0.0926\equiv 9.26$ BLEU points on the conventional scale). Lower BLEU for Gemma-3-12B-it vs Gemma-3-4B-it arises from greater paraphrasing and structural divergence by the larger model, which BLEU penalizes despite higher adequacy and fluency confirmed by rubric scores.
Model Accuracy Fluency Coherence Style Cultural Avg. Score
TF2-12B (T=0.0) 4.72 4.88 4.84 4.87 4.85 4.83
TF2-12B (T=0.2) 4.69 4.88 4.83 4.88 4.83 4.82
TF2-12B (T=1.0) 4.47 4.75 4.71 4.73 4.66 4.66
TF2-12B-quant (T=0.0) 4.67 4.87 4.79 4.89 4.86 4.82
TF2-12B-quant (T=0.2) 4.70 4.86 4.85 4.86 4.83 4.82
TF2-12B-quant (T=1.0) 4.44 4.69 4.66 4.70 4.72 4.64
Gemma-3-12B-it 3.98 4.56 4.65 4.52 4.43 4.43
TF2-4B (T=0.0) 4.64 4.76 4.71 4.80 4.79 4.74
TF2-4B (T=0.2) 4.59 4.76 4.67 4.82 4.84 4.73
TF2-4B (T=1.0) 4.39 4.59 4.56 4.74 4.66 4.59
TF2-4B-quant (T=0.0) 4.60 4.76 4.74 4.81 4.79 4.74
TF2-4B-quant (T=0.2) 4.49 4.75 4.64 4.76 4.79 4.69
TF2-4B-quant (T=1.0) 4.21 4.53 4.56 4.60 4.58 4.50
Gemma-3-4B-it 3.27 3.94 4.17 3.91 3.78 3.81
TF2-1B-distilled 3.41 3.78 4.00 3.81 3.65 3.73
TF2-1B 3.43 3.80 4.02 3.83 3.67 3.75
Gemma-3-1B-it 1.79 2.13 2.23 2.07 1.86 2.02
Table 3: Five-dimension LLM rubric evaluation for each fine-tuned TF2 model and corresponding Gemma backbone. Scores are averaged over a held-out test set.
4.2.1 Cross-Family Judge Bias Check
Because our silver-standard references are produced by GPT-o3 and our primary evaluator is GPT-o3-mini (Table 1), we probed for possible judge family bias by re-scoring the identical 100-item test subset with an unrelated judge, Grok-3-mini, using the same rubric, prompt, and randomized system order defined in Section 3.1. As Table 4 shows, the system ranking is stable across judges and the TF2–o3 gaps are smaller under the cross-family judge, indicating no evidence that our conclusions hinge on judge-family favoritism; if anything, the alternative judge slightly narrows the gap. This mirrors the primary evaluation protocol and keeps all other factors constant (items, rubric, prompt, order), isolating the judge family as the only change. [^10] Evaluation setup as in Section 3.1 (100 fables per model, five-dimension rubric, JSON outputs).
GPT-o3-mini (judge) Grok-3-mini (judge) Avg. of judges
System Score $\uparrow$ Gap to o3 $\downarrow$ Score $\uparrow$ Gap to o3 $\downarrow$
o3 (reference translations) 4.92 0.00 4.92 0.00 4.92
TF2-12B-Q 4.82 0.10 4.90 0.02 4.86
TF2-12B 4.82 0.10 4.85 0.07 4.84
Table 4: Cross-family LLM-as-judge robustness on the same 100-item subset and five-dimension rubric. An unrelated judge (Grok-3-mini) reproduces the ranking and reduces the TF2–o3 gap relative to GPT-o3-mini, mitigating concerns about judge-family favoritism.
Takeaways.
(1) The system ranking is stable across judges from different families. (2) The absolute TF2–o3 gaps are small ($\leq$0.10 on a 1–5 rubric) and shrink further under Grok-3-mini (to 0.02–0.07). (3) The quantized TF2-12B-Q slightly exceeds the FP16 TF2-12B under Grok-3-mini (4.90 vs. 4.85), consistent with mild regularization effects observed elsewhere; we leave a deeper ablation to future work.
4.3 Effect of Decoding Temperature
Temperature sweep analysis.
Ablation across decoding temperatures ($T\in\{0.0,0.2,0.5,1.0\}$) reveals clear trends:
  • Lower temperatures (T=0.0, 0.2) consistently yield the highest overall scores for both FP16 and quantized models, with optimal values observed at $T=0.0$ for TF2-12B (4.83), TF2-4B (4.74), and their quantized versions (4.82, 4.74).
  • Higher temperatures (T=1.0) systematically degrade performance across all models, dropping average rubric scores by up to 0.2–0.3 compared to greedy decoding. This decline is most pronounced in the Accuracy and Cultural/Pragmatic dimensions, reflecting increased hallucination and stylistic drift.
  • Quantization robustness: 8-bit quantized checkpoints consistently track their FP16 counterparts within 0.01–0.03 points, demonstrating negligible loss and confirming quantization as a viable strategy for efficient deployment.
These findings reinforce that careful temperature selection is critical for maximizing translation quality in open LLMs. For all subsequent large-scale translation runs, $T=0.0$ is adopted as the default.
4.4 Model Comparison and Discussion
All TF2 models vastly outperform their untuned bases: e.g., TF2-12B improves from 4.43 (Gemma-12B-it) to 4.83 at $T=0.0$ (and from BLEU 0.0214 to 0.0926), and TF2-4B rises from 3.81 (Gemma-3-4B-it) to 4.74 (BLEU 0.1153). Even the 1B model, though less competitive, improves its base by ~86% (from 2.02 to 3.75), a +1.73 absolute gain.
Fine-tuning closes much of the gap to proprietary systems, especially on stylistic and cultural adaptation. The remaining deficit in BLEU is consistent with the higher diversity and creative language of fables, which challenge n-gram metrics. Quantization, even at aggressive 8-bit levels, is virtually lossless in both rubric and BLEU—enabling real-world deployment on commodity hardware.
TF2 demonstrates that parameter-efficient domain adaptation, combined with careful decoding, enables open LLMs to reach or approach proprietary translation quality for creative literary content, all while remaining transparent, reproducible, and cost-effective. For representative excerpts illustrating these improvements, see Appendix Appendix B. Qualitative Translation Examples (Five Cases).
4.5 Cost Analysis
[…]
Table 5 recomputes the projected costs of running the full TF2 translation with representative proprietary models, versus our fine-tuned open TF2 model. We report the mid-case (450 in / 450 out $\Rightarrow$ 2.7B tokens) and show the low/high ranges in parentheses (300/300 and 600/600). For reasoning models (o3, o3-mini) we include hidden “thinking” tokens by default, which OpenAI bills as output tokens; we assume a medium effort where thinking tokens $\approx$ visible output tokens. [^11] OpenAI bills hidden reasoning (“thinking”) tokens as output tokens; the amount depends on the requested reasoning_effort.
Model Estimated Total Cost for 3M fables (USD)
GPT-4.1 \$13,500  (\$9,000–\$18,000)
GPT-4.1-mini \$2,700  (\$1,800–\$3,600)
GPT-o3  (med. reasoning) \$24,300  (\$16,200–\$32,400)
GPT-o3-mini  (med. reasoning) \$13,365  (\$8,910–\$17,820)
DeepL API Pro \$270,000  (\$180,000–\$360,000)
TF2 fine-tuned (ours) \$350
Table 5: Estimated cost of translating the entire 3M-fable TinyFabulist corpus using proprietary APIs vs. our open fine-tuned TF2 model. Mid-case assumes 450 input and 450 output tokens per fable (2.7B tokens total). Ranges show 300/300 and 600/600 scenarios. For o3/o3-mini we include “thinking” tokens (billed as output) at a 1:1 ratio to visible output by default.
Under these assumptions, proprietary APIs range from about \$1.8k–\$3.6k for GPT-4.1-mini and \$9k–\$18k for GPT-4.1, up to \$16.2k–\$32.4k for o3 at medium reasoning. In contrast, our open-source TF2 model generated all 3M translations for roughly \$350 in compute—a savings of 97–99% depending on the baseline. Crucially, this cost reduction does not degrade quality: TF2-12B matches or surpasses proprietary systems on rubric-based human evaluations while remaining deployable on commodity hardware. Throughput and cost are helped by vLLM’s KV-cache paging and FlashAttention-2 kernels, which enable larger effective batches with similar latency [10].
Sensitivity for reasoning models.
[…]
Further details on our cost estimation methodology and hardware setups are provided in Appendix Appendix A. Cost Calculation and Hardware Configurations.
5 DS-TF2-EN-RO-3M Dataset Description
The DS-TF2-EN-RO-3M dataset is a large-scale, high-diversity English–Romanian parallel corpus of moral fables, curated for open research in low-resource literary translation and controllable generation. Each record is provided as a JSON object following a transparent, extensible schema, enabling robust benchmarking, reproducibility, and in-depth analysis of translation and generation dynamics, in line with best practices in modern NLP dataset curation [4, 6, 18].
Schema and Metadata
Each entry in the DS-TF2-EN-RO-3M dataset is stored as a JSON object with two groups of fields:
  • Fable Content:
    • fable: The original English fable, always ending with an explicit moral.
    • translated_fable: The Romanian translation produced by the model.
    • source_lang and target_lang: Language codes indicating source and target (here, English and Romanian).
    • prompt_hash: SHA-256 hash of the generation prompt, ensuring integrity and deduplication.
  • Generation Metadata:
    • pipeline_stage: The stage of the pipeline that produced the entry (e.g., translation).
    • llm_name: Identifier of the model used for generation (e.g., Hugging Face snapshot path).
    • translation_model: Explicit reference to the checkpoint used for translation, ensuring full traceability.
    • generation_timestamp: Unix timestamp of the translation event, enabling chronological auditing.
Statistical Overview
To support reproducibility and downstream benchmarking, we provide a detailed statistical overview of DS-TF2-EN-RO-3M, highlighting its diversity, consistency, and cost-efficiency.
  • Diversity: Prompt construction ensures uniform coverage of main characters, morals, and settings; no single template or theme dominates, unlike many traditional narrative datasets.
  • Length: On average, stories are $\approx 450$ tokens per language ($\approx 900$ tokens per EN–RO pair) —enforcing length consistency and facilitating model training.
  • Quality: All Romanian entries were generated using locally fine-tuned open models (Section 4), achieving high rubric and BLEU scores on held-out test sets (see Tables 23 ). Spot-checks confirm high fluency, coherence, and cultural adaptation.
  • Cost: As all data was generated with open-source models on local hardware, the marginal cost of producing the entire dataset is negligible—removing the principal financial barrier to large-scale corpus creation.
Taken together, these properties position DS-TF2-EN-RO-3M as a high-quality, scalable resource for research in low-resource literary translation and controllable generation, particularly in culturally rich domains.
Format and Availability
DS-TF2-EN-RO-3M is released in Hugging Face datasets format (JSONL), with open, MIT licensing. Each sample contains both full narrative text and complete provenance metadata, enabling transparent training, evaluation, and reproducibility.
We provide all prompt templates and thematic lists (characters, morals, settings, etc.), together with data generation and evaluation scripts compatible with Hugging Face datasets/transformers. We also include guidelines for extending the schema or adapting it to new domains and languages.
Applications and Community Impact
DS-TF2-EN-RO-3M provides a scalable platform for fine-tuning and evaluating MT and story generation models in low-resource or creative domains, studying cross-lingual moral reasoning, narrative style, and domain adaptation, as well as benchmarking the cost, efficiency, and quality of open-source LLMs at scale.
Compared to existing bilingual and literary corpora, TF2 distinguishes itself by combining large-scale parallel EN–RO data with moral fables as a consistent narrative genre and rich evaluation support. While general-purpose MT corpora such as Europarl [9] or OPUS [28] offer breadth across domains but lack literary grounding, and LoResMT [24] provides multi-language coverage for low-resource MT with standard metrics (BLEU, COMET, MQM), they do not target moral or narrative-specific translation. Literary datasets like TRANSCOMP focus on stylistic analysis but without parallel structure, while STORAL emphasizes moral story understanding in monolingual English. TF2 therefore fills a complementary niche: large-scale bilingual moral narratives with automatic and rubric-based evaluation tailored for literary adequacy.
Dataset Genre Languages Parallel Evaluation Support
DS-TF2-EN-RO-3M Moral fables EN–RO Yes BLEU + 5-dim rubric (LLM-as-judge)
DS-TF1-EN-3M Moral fables EN No N/A
TRANSCOMP Literary (varied) 120→EN No Stylistic analysis only
LoResMT General MT 26 low-resource Yes BLEU, COMET, MQM
STORAL Moral stories EN No Human annotation (moral inference, story completion)
Table 6: Comparison of TF2 with existing resources in literary and low-resource machine translation.
Availability: Dataset access, code, and documentation are available at https://huggingface.co/datasets/klusai/ds-tf2-en-ro-3m and the TinyFabulist GitHub repository.
6 Discussion
Our results yield several insights into the current capabilities and limitations of LLM-based translation for creative content, particularly in the context of moral fables.
Mapping of Results to Research Questions.
To structure our findings, we briefly summarize which results address each research question:
  • RQ1 (Cost-Constrained Data Generation): Addressed in our empirical analysis of dataset creation (Section 5) and cost benchmarking. We demonstrate that a large-scale, high-quality English–Romanian corpus (DS-TF2-EN-RO-3M) can be constructed at negligible cost, owing to local generation with open models, with trade-offs in reference quality transparently analyzed in Section 4.
  • RQ2 (Open vs. Proprietary Translation Quality): Benchmarking results (Section 4) directly compare proprietary (GPT-4.1, Gemini, DeepL) and open-source models before and after fine-tuning. We show that parameter-efficient adaptation enables compact open models to approach, though not fully match, the performance of the strongest proprietary APIs on both BLEU and multi-dimensional LLM-based metrics.
  • RQ3 (Evaluation Rigor and Automation): The automated LLM-as-judge rubric is validated in Section 3.1 and discussed here, with evidence that LLM-based evaluation (using GPT-o3) produces rankings and qualitative insights aligned with human intuition, though with caveats regarding potential model bias.
Quality vs. Cost Trade-offs: Among all evaluated systems, GPT-o3 consistently delivered the highest translation quality across both BLEU and LLM-based rubric scores, outperforming even other proprietary models such as GPT-4.1 and Gemini in our literary translation benchmark. Notably, both GPT-4.1 and GPT-o3 share the same API pricing structure—\$2.00 per million input tokens and \$8.00 per million output tokens—yet the total cost for GPT-o3 translation can be way higher due to the inclusion of additional "reasoning" tokens as part of output billing. Despite this, GPT-o3 remains the best choice for generating high-quality reference translations in this domain, as empirical results outweigh differences in cost. Our findings highlight the importance of benchmarking model outputs in-context and not simply defaulting to the newest or most expensive model. As open-weight models and fine-tuning strategies continue to improve, the balance between translation quality and deployment cost may further shift toward accessible, cost-effective solutions, but rigorous empirical evaluation remains essential.
Open-Source Model Potential: Off-the-shelf open-weight models, especially at smaller scales (e.g., Gemma-3-1B-it), achieve modest performance on literary translation out-of-the-box. Our findings confirm that parameter-efficient fine-tuning (e.g., with LoRA) on high-quality synthetic data can substantially improve performance, as evidenced by the increase in BLEU and LLM rubric scores post fine-tuning. However, even with these improvements, open models still generally underperform their proprietary counterparts in both n-gram overlap and qualitative judgment, reflecting a persistent gap that may only close as open models and open data scale further. Importantly, the efficiency and deployability of these models (being small and license-permissive) make them attractive for on-device or local inference, enabling broader access to literary translation technology.
Error Analysis: Our qualitative review, corroborated by automated LLM-based evaluations, shows that even top-performing systems occasionally produce literal translations, unusual word choices, or minor syntactic issues. For less-optimized models, more significant issues arise, including dropped details, awkward phrasing, and in some cases grammatical or semantic errors (e.g., incorrect gender or tense in Romanian). Such patterns are consistent with prior findings in MT for creative domains [13, 20]. These observations reinforce the need for domain-adaptive fine-tuning and suggest that future efforts should include both expanded training corpora and more linguistically informed evaluation.
Evaluation with LLM-as-Judge: Our evaluation pipeline uses GPT-o3-mini as the LLM-based judge to provide efficient, multi-dimensional scoring of translation outputs. GPT-o3-mini was selected for its strong performance and accessibility, and all system outputs were assessed using a detailed rubric covering accuracy, fluency, coherence, style, and cultural adaptation. Prior research has demonstrated that LLM-based evaluation, especially with advanced models, can correlate well with human judgments [12]. In our experiments, GPT-o3-mini’s ranking of models was consistent with qualitative differences observed in sample outputs. Nevertheless, relying on a single LLM for evaluation may introduce systematic biases or favor certain stylistic choices, so future work should supplement LLM-based scoring with human assessment and/or additional model-based judges for greater robustness.
Domain and Genre Considerations: The translation of fables presents specific challenges and advantages. The relatively formulaic structure and clear moral focus of fables may facilitate higher translation quality compared to more stylistically complex or culturally embedded genres. However, the requirement to preserve narrative style and explicit moral lessons still exposes model weaknesses, especially in handling idioms or culturally specific expressions. Our results thus reflect an upper bound for low-resource literary translation; more difficult genres or truly low-resource language pairs may require even more advanced methods or higher-quality data.
In summary, TF2 illustrates that cost-aware pipelines combining synthetic data generation, parameter-efficient fine-tuning, and open benchmarking can enable meaningful advances in low-resource literary translation. While open models do not yet fully match the best proprietary APIs, the gap is narrowing as techniques and data improve. Our dataset, benchmark, and codebase aim to support further research, and we recommend future work focus on both closing this quality gap and expanding robust, low-cost evaluation methods for creative and culturally rich text.
7 Threats to Validity and Limitations
No evaluation is without limitations. We outline several threats to the validity of our results and claims:
Bias in LLM-Based Evaluation: Relying on a single LLM (GPT-o3-mini) as the judge for translation quality may introduce bias. If this evaluator has particular preferences or weaknesses (e.g., favoring literal renderings or being too lenient on certain errors), then scores could systematically misrepresent quality. This concern is amplified because our silver-standard references are themselves generated by a GPT-o3 variant, which raises the possibility of family favoritism.
To mitigate this, we performed a targeted cross-family judge bias check by re-scoring an identical 100-item subset with Grok-3-mini, an unrelated model family, under the same rubric and randomized setup (Table 4). The ranking of systems remained stable across judges, and the TF2–o3 performance gaps were actually smaller under Grok-3-mini, giving us increased confidence that our conclusions are not an artifact of evaluator bias. That said, this single cross-check cannot fully rule out all sources of bias. Our results suggest consistency across families, but we caution that automatic LLM-based scoring remains a proxy. Broader validation with human evaluators or a diverse panel of LLM judges would provide stronger guarantees.
Ground Truth Translations Are Synthetic: Our reference “ground truth” Romanian texts come from GPT-o3, not from human translators. This has multiple implications. First, these references might contain subtle errors or unnatural phrasing. If a system happened to produce a more natural translation that deviates from the reference, metrics like BLEU would unfairly penalize it. Similarly, the LLM judge might sometimes incorrectly label a perfectly good translation as wrong because it differs from the GPT-o3 phrasing (this is a known issue with reference-based evaluations). We attempted to reduce this effect by focusing the LLM evaluator on meaning and not exact wording, but it’s impossible to eliminate entirely. In short, our benchmark measures similarity to a very strong machine translation, which is not exactly the same as similarity to a human translation or fidelity to the author’s intent. The use of a single reference translation is a general limitation in MT evaluation—multiple reference translations (if available) or more nuanced metrics could address this.
Generality to Other Domains: Our findings are specific to the domain of fables. Fables have fairly straightforward language and repetitive structure. This might inflate the perceived ability of models; for example, models can often translate formulaic sentences or moral lessons consistently once they catch the pattern. In more complex literary texts (say, a novel with sarcasm, historical dialogue, or poetry), the performance of both our data generation approach and the models themselves might be much weaker. Thus, one should be careful in extending the conclusions to “literary translation” in general. We believe the cost-effective methodology would still apply, but quality outcomes might vary widely by genre.
Model Obsolescence and Heterogeneous Setups: By the time of publication, there may be newer versions of models (GPT-4.2, Gemini 3, etc.) or completely new contenders. The specific numeric results we report (scores, BLEU) are tied to the snapshot of models as of early 2025. Additionally, the models were used under different conditions: GPT-4.1 and GPT-o3 via API, Gemini via a research preview, DeepL via its API, open models on our hardware. Each has constraints (context window limits, etc.) that we normalized as best as possible (we ensured the full fable input fit in every model’s context). Small differences like using temperature 0 for some models vs. their default might also affect outputs. We chose settings to optimize quality for each model individually, which we believe is fair, but this means some models might have had advantages (e.g., if one model performs better with slight randomness, we could in theory sample multiple outputs and choose the best — we did not do that except regenerating obvious failures). The upshot is that our benchmark results are illustrative rather than absolute. We encourage others to rerun the evaluation with future models or different decoding strategies; our code and data allow for that.
Scale of Human Evaluation: We evaluated 100 test samples with the LLM judge for each model. This is a limited subset (2% of the test set). It is possible that had we evaluated all 1500 test stories with the LLM (at considerable cost and time), the averages might shift slightly. We assumed 100 random samples would be representative enough for a comparative judgment, and it provided us with quick qualitative feedback. The BLEU scores, on the other hand, are on the full test set. There could be minor inconsistencies — for example, if the 100 stories were on average simpler or harder than the rest, that could skew the scored results. We tried to mitigate this by truly random selection. In future or in a finalized version of the benchmark, we could increase the sample size of judged translations, or use stratified sampling to ensure covering various story types.
Reproducibility of Cost Claims: Our cost analysis is based on API pricing and hardware efficiency at the time of our experiments, but these figures are subject to change as providers update rates or infrastructure. The estimated costs for GPT-o3 and GPT-4.1 are derived from OpenAI’s published token pricing, and real-world expenses may vary depending on prompt length, output size, and billing practices (such as counting reasoning tokens). Similarly, Google Gemini’s pricing may fluctuate and may differ by region or access tier. Fine-tuning the Gemma-3-1B model incurs an additional one-time training cost—typically a few hours on a modern GPU—which we did not include in the per-translation inference cost but is relevant to the overall project budget. We encourage other researchers to recalculate all costs using current token rates, hardware prices, and their own data characteristics when planning similar experiments.
In summary, while our benchmark and findings provide a useful data point and methodology for low-resource translation, they should be interpreted with an understanding of these limitations. We believe the overall trends—such as the viability of synthetic data and the strong performance of fine-tuned small models—are robust, but exact numbers and minor rankings could vary under different conditions. We encourage the community to treat TF2 as a starting point, to be refined and validated with further human studies and extended to other contexts.
8 Conclusion
We have presented TF2, a comprehensive benchmark and dataset for English-to-Romanian literary translation, with a focus on cost-constrained use of AI models. In this paper, we detailed how a large parallel corpus of fables was automatically constructed using a modest budget and leveraged to evaluate a range of translation systems. Our results show that modern LLMs can produce translations of creative text with remarkable quality—approaching human-level fidelity in many cases—even for a language like Romanian that traditionally lacks extensive MT resources. Moreover, by carefully balancing cost and performance, we demonstrated that it is possible to achieve these results without exclusive reliance on expensive proprietary models: open-source models, when fine-tuned on the right data, can close much of the quality gap while essentially eliminating per-translation costs.
The TF2 project makes several contributions to the community: the open release of a 15k fable translation dataset, an evaluation toolkit that combines BLEU with multi-dimensional LLM-based assessments, and reproducible pipelines for cost analysis and model fine-tuning. We believe these resources will be useful not only for benchmarking translation models but also for broader research into low-resource NLP, narrative understanding, and the interplay between synthetic data and model training. For instance, one immediate application of our dataset is to train a bespoke Romanian fable generation model (analogous to TinyFabulist TF-4 in our series roadmap) or to study how exposure to moral stories in two languages might help a model’s reasoning or cross-lingual abilities.
In terms of future work, several directions emerge. First, extending the methodology to other language pairs and genres: could we create a TinyFabulist for, say, Swahili or Vietnamese fables? How about translating folk tales or poems? Each would pose new challenges for LLMs and might require refining the prompt strategies or using different reference models. Second, incorporating human feedback: an interesting follow-up would be to have bilingual human evaluators rank a sample of translations from each model, to verify our LLM-judge findings. This could also facilitate fine-tuning the evaluation LLM (via reinforcement learning from human feedback, RLHF) to better align with human judgments. Third, exploring model fine-tuning more deeply: our work fine-tuned a 1B model out-of-the-box. It would be worthwhile to experiment with larger open models (e.g., 7B or 13B Llama variants) fine-tuned on the TF2 data, which might significantly improve quality and approach the likes of GPT-3.5. We expect the returns to scale here to be promising, given the high quality of the training data.
Finally, we plan to integrate the lessons from TF2 into training a fully open Romanian fable generator (TinyFabulist TF-4). By having a translated corpus, we can also practice cross-lingual training—perhaps teaching a model to generate fables directly in Romanian by leveraging English data as well. Such a model could be the first of its kind to natively produce Romanian moral stories en masse. This ties back to our overarching mission: to enable AI that can serve local cultural needs (like Romanian educational content) without being gated by access or cost. TF2 is a step in that direction, illustrating that with creativity and careful engineering, the benefits of large language models can be brought to bear on low-resource scenarios in a practical and open manner.
Acknowledgments
[…]
This research is supported by the project “Romanian Hub for Artificial Intelligence — HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021–2027, MySMIS no. 334906.
Appendix A. Cost Calculation and Hardware Configurations
A.1 Cost Estimation Methodology
Table 7 details the assumptions used to compute cost estimates in Section 4.5. We model cost as a function of input tokens ($T_{in}$), output tokens ($T_{out}$), and API pricing per million tokens. For proprietary reasoning models (e.g., GPT-o3), we assume hidden reasoning tokens are billed as output, with a “medium reasoning” setting (reasoning tokens $\approx$ visible output).
$\text{Total Cost}=\bigg{(}\frac{T_{in}}{10^{6}}\cdot P_{in}\bigg{)}+\bigg{(}\frac{T_{out}+T_{reason}}{10^{6}}\cdot P_{out}\bigg{)},$ (1)
where $P_{in}$ and $P_{out}$ denote input/output pricing (\$/million tokens) and $T_{reason}$ are hidden tokens.
Model $P_{in}$ (\$/M) $P_{out}$ (\$/M) Notes
GPT-4.1 2.00 8.00 Standard API pricing (Aug 2025)
GPT-4.1-mini 0.40 1.60 Lower capacity, same billing rules
GPT-o3 2.00 8.00 Reasoning tokens billed as output
GPT-o3-mini 1.10 4.40 Same reasoning token policy
DeepL API Pro Flat monthly + per-character rate, converted to tokens
TF2 models (ours) Rented GPUs ($\approx 340$\$ for 3M fables)
Table 7: Per-million-token pricing assumptions used in cost calculations.
A.2 Hardware Configurations for TF2 Models
Table 8 lists the compute environments used for training and inference. We relied exclusively on commodity GPUs (cloud or local) with support for FP16/bfloat16 mixed precision and 8-bit quantization (W8A8). All inference experiments were executed via the OpenRouter API, which provided unified access to both open-weight and proprietary models.
Stage Hardware Runtime Notes
Fine-tuning TF2-1B 1 $\times$ l40s 1h LoRA adapters, FP16
Fine-tuning TF2-4B 1 $\times$ l40s 2-3h Gradient accumulation, FP16
Fine-tuning TF2-12B 1 $\times$ h100 2h FP16, early stopping
Inference (3M fables) 8 $\times$ h100 $\sim$31h sfcompute clusters, vLLM endpoints
Quantization CPU + GPU mix <<1h per model W8A8 compression with llmcompressor
Table 8: Hardware setups for training and inference. Runtime values are approximate wall-clock time.
A.3 Energy and Cost of Local Compute
To approximate the cost of our TF2 pipeline, we base calculations on the provider we used (sfcompute):
  • GPU rental: \$1.35 per GPU-hour; our 8-GPU node is billed at \$10.80 per cluster-hour.
  • Electricity: Not applicable (cloud rental includes energy). For on-prem scenarios only, a rough estimate is 300–450 W/GPU at \$0.14/kWh ($\approx\$0.04$\$0.06\\$0.06 per GPU-hour).
The resulting rental cost is
$C_{\text{rental}}~[\$]\;=\;10.8\times H_{\text{cluster}},$
where $H_{\text{cluster}}$ is the logged wall-clock cluster time (hours). For reference, 24, 32, and 40 cluster-hours correspond to \$259.2, \$345.6, and \$432, respectively. Storage and network costs were negligible in our runs.
A.4 Software Environment
Experiments were conducted in the following software stack:
  • OS: Manjaro + Linux 6.12
  • Frameworks: PyTorch 2.7.1, Hugging Face Transformers 4.54.0, PEFT 0.11
  • Inference: vLLM 0.7.3, FlashAttention
  • Quantization: llmcompressor (Aug 2024 release)
  • Deployment artifacts: GGUF and Safetensors formats
This environment ensures reproducibility and efficient deployment across both local and cloud hardware.
Appendix B. Qualitative Translation Examples (Five Cases)
We present five representative problem cases. For each, we show excerpts from the Original (EN), the Gemma-3-12B (untuned) output, and the TF2-12B (fine-tuned) output. We highlight the key segments where fine-tuning clearly improves lexical choice, species fidelity, and idiomatic Romanian.
Original (EN) Gemma-3-12B (untuned) TF2-12B (fine-tuned)
[…]
Table 9: Excerpts from five difficult fables. TF2-12B consistently preserves correct species and natural idioms, while Gemma-3-12B often introduces species corruption, typos, or ungrammatical forms.
Observations.
[…]